Pat travels 70 miles on her milk route, and Bob travels 75 miles on his route. Pat travels 5 miles per hour slower than Bob, and her route takes her one half hour longer than Bob's. How fast is each one travelling?
Answer by dkppathak(439)

Pat travels 70 miles on her milk route, and Bob travels 75 miles on his route. Pat travels 5 miles per hour slower than Bob, and her route takes her one half hour longer than Bob's. How fast is each one travelling?
let bob travel X miles/hours and pat X-5 miles /hours
T1=75/x
t2=70/x-5
t2-t1=1/2
70/x-5-75/x=1/2
70x-75(x-5)/x^2-5x=1/2
(-5x+375)x2=x^2-5x
-10x+750=X^2-5x
x^2+5x-750=0
x=-30 or x=25 rejecting negative value x=25 miles/hours